OpenStudy (anonymous):

Find the slope and y intercept 7-1/2y=2x

OpenStudy (shadowfiend):

Do you know of a form of equations where you can easily get the slope?

OpenStudy (anonymous):

do you mean y=mx+b

OpenStudy (shadowfiend):

RIght.

OpenStudy (shadowfiend):

So if we can solve the above equation for y, then we might be closer to getting the slope and y-intercept, right? Can you solve the equation for y?

OpenStudy (anonymous):

Well I do know that you take 7 and subtract that from both sides geting -1/2y=2x-7 right?

OpenStudy (shadowfiend):

Exactly, that's the first step.

OpenStudy (anonymous):

okay then after that im confused

OpenStudy (anonymous):

well 2 would be the slope right?

OpenStudy (shadowfiend):

So then you need to get the y alone. That means you have to get rid of the -1/2 that's next to it.

OpenStudy (anonymous):

okay so 1/2+ -7 = 3.5

OpenStudy (shadowfiend):

Close. Because y is multiplied by the -1/2, we have to divide both sides by -1/2.

OpenStudy (shadowfiend):

Dividing by -1/2 is the same as multiplying by -2.

OpenStudy (anonymous):

oh thats where i always mess up

OpenStudy (shadowfiend):

So we would have: -2(-1/2y) = -2(2x - 7)

OpenStudy (anonymous):

-4x+14

OpenStudy (shadowfiend):

Exactly :)

OpenStudy (shadowfiend):

So now what's the slope?

OpenStudy (anonymous):

YAY haha the slope would be -4

OpenStudy (shadowfiend):

Right! And the y-intercept?

OpenStudy (anonymous):

14

OpenStudy (shadowfiend):

Rock it :)

OpenStudy (shadowfiend):

Ok, so now let's look at your other question and see if we can apply the same idea.

OpenStudy (anonymous):

okay sounds good

OpenStudy (anonymous):

(-2,8); 7x=5y+6

OpenStudy (shadowfiend):

Well, hehe, I was going to actually go to that problem, so we keep things organized :)

OpenStudy (anonymous):

o okay ill head over to it sorry